OpenStudy (anonymous):

Show that 1) square root of i = 1+i/square root of 2 2) square root of -i = 1-i/square root of 2 Please help me!! Thanks you all! :-)

OpenStudy (harsimran_hs4):

lets start step by step 1. assume the square root to be a + ib the (a + ib)^2 = i now compare the real and imaginary parts and tell me what you get

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

I have no idea! :P

OpenStudy (harsimran_hs4):

ok (a + ib)^2 = a^2 + (ib) ^ 2 + 2a(ib) = a^2 -b^2 + 2abi = i so comparing we get a^2 - b^2 = 0 and 2ab = 1 now can you solve for a and b?

OpenStudy (harsimran_hs4):

did you manage to get hold of something?